OpenStudy (anonymous):

Robbie the Robot is on a weather satellite orbiting Earth about 3600 km above the surface. The Earth’s radius is about 6400 km. He has had a malfunction in his output device, and the satellite is traveling without communication. His last report was only in terms of trigonometric values and was only partially received. It said, sin Θ < 0… and then he was lost again. Part 1: Create a set of coordinates that would be reasonable for Robbie’s position in space and satisfy his last, partial report. Using complete sentences, describe Robbie’s location and your reasoning.

OpenStudy (anonymous):

Is this science ir math

OpenStudy (geekfromthefutur):

It asks for the values of the Sin(theta),Cos(theta) and Tan(theta) using your coordinate point. So basically this is what you do, you take, For Sin its, Sin=1736 / 10000 = 0.1736 . that's your Sin(theta) values. Then for Cos(theta its, Cos=9848/10000= 0.9848 that's your cos(theta) v for your last part which is, tan, its tan=1736 / 9848= 0.1763 So you have, Sin=1736/1000=(0.1736) Cos=9848/1000=(0.9848) Tan=1736/9848=(0.1763)

OpenStudy (anonymous):

how did you get the values for the point though? @geekfromthefutur

OpenStudy (geekfromthefutur):

What do you mean